OpenStudy (anonymous):

Write the decimal 0.082 as a fraction in simplest form.

OpenStudy (callisto):

0.1 = 1/10 0.01 = 1/100 0.001 = 1/1000 0.11 = 11/100 0.111= 111/1000 0.082 = ...?

OpenStudy (anonymous):

x.x blehh lost